Ques1) A caterpiller of length 4cm crawls in the direction of electron drift along a 5.2mm diameter bare copper wire that carries an uniform current of 12 A. (a) What is the potential difference b/w the two ends of the caterpiller?  (b) Is it's tail positive or negative relative to the head? (c) How much time does it takes to crawl 1 cm if it crawlsat the drift speed of the electrons in the wire? (The number of vharge carries per unit volume is 8.49 x 10 28 m-3.)


Ques2) What would happen to the drift velocity of the electrons in a wire and to the current in the wire if the electrons could move freely without the resistance through the wire?

Grade:12

1 Answers

Ramesh V
14 years ago

(1)

Area of copper wire : Pi/d*d2 = 2.123x10-5 m2 

lenght of catrepillar : L : 0.04 m

Resistivity of copper is: 1.69 x 10-8 ohm-meter

no. density of charge carriers in copper: ρ = 8.49 x 10 28 m-3

 (a) volatge V = iR = iρL/A

                 V = (12*1.69 * 10-8 * 4*10-2) / 2.123x10-5

                    = 0.382 mV

  (b) current i is opposite to electron drift

        i runs from Vhig to Vlow

        So, the head is at higher potential.

 (c)  Vd = drift velocity

        Δt  = Δx / Vd

        Vd = J / ne = i / Ane

       Δt  = Δx .Ane /i

             = 10-2 * 2.123x10-5 *8.49*1028 *1.602*10-19 / 12

             =  240.5 sec

time it takes to crawl is 240.5 sec

----

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch